# Graph Convolutions Enrich the Self-Attention in Transformers!

### v2

@@ 10  VIOsPAPR_RTCE *rtce;
Appendix K Code Clone Detection
K.1 Detailed Experimental Settings
@@
Code clone detection aims to measure the similarity between two code snippets and predict whether they have the same functionality. We experiment with the Java data provided by Wang et al. (2020).
@@
We build our experiments on top of the open-sourced code  [^8] https://github.com/salesforce/CodeT5 and recipes provided by Wang et al. (2021b).
@@
Appendix L Time Complexity and Empirical Runtime Analysis
@@
We report the training time of various methods with GFSA in Tables 31 to 23. In general, the training time of methods with GFSA is slightly longer than that of existing methods. For example, the Transformer for the automatic speech recognition task increases from 190.5 seconds to 191.6 seconds on Librispeech 100h dataset, as increases of only 1 second. Instead of computing higher-order polynomial terms, our GFSA approximates them, with only a small increase in runtime, which is not very significant.
@@
Table 21: Training time (seconds per epoch) on graph classficiation tasks
@@
Appendix M Inference Time
We report the inference time of various methods with GFSA in Tables 24 to 29.
@@
Table 27: Inference time on graph classficiation tasks
@@
Appendix N Frequency Analyses in the Singular Value Domain
Graph signal processing (GSP) (Sandryhaila & Moura, 2013; 2014) can be understood as a generalized concept of DSP — in other words, DSP is a special case of GSP where a line graph with $n$ nodes is used and therefore, the graph Fourier transform (GFT) of the line graph is identical to the discrete Fourier transform.
However, one can use the singular value decomposition, $\bm{S}=\bm{V}^{\intercal}\bm{\Lambda}\bm{U}$, when the GSO is not diagonalizable but symmetrically normalized instead of the eigendecomposition (Maskey et al., 2023). Both the singular value decomposition and the eigendecomposition project the original signal onto a set of basis, but they use different basis sets. In the singular value decomposition, we sort the set of basis in ascending order of their eigenvalues, i.e., $\bm{\Lambda}$, and perform frequency domain-like analyses (Zou et al., 2022; Maskey et al., 2023).
Since the self-attention matrix’s row-wise sum is always 1, the following is the case: $\bar{\bm{A}}=\bm{D}^{-1}\bm{A}=\frac{1}{n}\bm{A},$ where $n$ is the number of tokens. Maskey et al. (2023) define the following symmetrically normalized adjacency matrix (SNA): $\bm{D}_{in}^{-1/2}\bm{A}\bm{D}_{out}^{-1/2}.$ Since the degree of every node is $n$ in the self-attention matrix, the following is the case: $\bm{D}_{in}^{-1/2}\bm{A}\bm{D}_{out}^{-1/2}=\bm{D}^{-1/2}\bm{A}\bm{D}^{-1/2}=\frac{1}{\sqrt{n}}\bm{A}\frac{1}{\sqrt{n}}=\frac{1}{n}\bm{A}=\bar{\bm{A}}.$ Therefore, the self-attention matrix is a special case of SNAs.
Appendix O Matrix Polynomial vs. Graph Fourier Transform
There are two paradigms of implementing graph filters: i) matrix polynomial, which does not require diagonalizability, and ii) graph Fourier transform, which uses the eigendecomposition for diagonalizable adjacency matrices or the Jordan decomposition for non-diagonalizable adjacency matrices.
Those two paradigms have their own weaknesses: i) the matrix polynomial approach requires explicit matrix multiplications, and ii) the graph Fourier transform approach requires expansive spectral decompositions. The matrix polynomial is preferred when the number of matrix multiplications is not many. Otherwise, the graph Fourier transform approach may be better since the matrix multiplication can be simplified after the eigendecomposition or the Jordan decomposition or the Jordan normal form.
Appendix P Approximation of $\bar{A}^{K}$ in  Equation 4
The general formulation of first-order taylor approximation at point $a$ is as follows:
$f(x)\simeq f(a)+f^{\prime}(a)(x-a).$ (16)
We approximate $f(K)=\bar{\bm{A}}^{K}$ with first-order taylor approximation at point $a=1$:
$f(K)=\bar{\bm{A}}^{K}\simeq f(1)+f^{\prime}(1)(K-1).$ (17)
Inspired by Brouwer et al. (2019), which approximates the derivative of hidden states of RNNs as the difference between hidden states, we effectively approximate $f^{\prime}(1)$, the derivative of $\bar{\bm{A}}^{K}$ estimated at the position of $K=1$, with the difference term as $f(2)-f(1)=\bar{\bm{A}}^{2}-\bar{\bm{A}}^{1}$. Then the approximated $\bar{\bm{A}}^{K}$ becomes as follows:
$f(K)=\bar{\bm{A}}^{K}\simeq f(1)+(\bar{\bm{A}}^{2}-\bar{\bm{A}})(K-1)=\bar{\bm{A}}+(K-1)(\bar{\bm{A}}^{2}-\bar{\bm{A}}).$ (18)
This approximation has two advantages:
  1. 1.
    The approximation for $\bar{\bm{A}}^{K}$ with $\bar{\bm{A}}$ and $\bar{\bm{A}}^{2}$ provides a simpler computation that can significantly reduce the required computational resources and time (see theoretical analysis in Section 3.2).
  2. 2.
    When approximating derivative of $\bar{\bm{A}}^{K}$ with difference term for graph filter design, our filter can reduce to GPR-GNN and GREAD, two popular GNN methods preventing the oversmoothing for graph convolutional networks, which are special cases of our design (see discussion in Section 3.3). In other words, our method is more general than those two GNNs.
Appendix Q GFSA in Selected Layers: A Strategy for Model Efficiency
Because GFSA requires more calculation than the original self-attention, the runtime after using GFSA increases. While the experiments in the main paper used GFSA on all layers of Transformer, we report a strategy to alleviate the computational burden of the Transformer models by applying GFSA only to some layers. For this purpose, GFSA is used only on even-numbered layers. In Tables 30 to 34, the results show that this strategy can reduce the increase in runtime and maintain performance compared to using GFSA for all layers.

### v5

@@ 10 VIOsPAPR_RTCE *rtce;
Appendix Q Code Clone Detection
Q.1 Detailed Experimental Settings
@@
Code clone detection aims to measure the similarity between two code snippets and predict whether they have the same functionality. We experiment with the Java data provided by Wang et al. [83].
@@
We build our experiments on top of the open-sourced code  [^9] https://github.com/salesforce/CodeT5 and recipes provided by Wang et al. [84].
@@
Appendix R Time Complexity and Empirical Runtime Analysis
@@
We report the training time of various methods with GFSA in Tables 23 to 28. In general, the training time of methods with GFSA is slightly longer than that of existing methods. For example, the Transformer for the automatic speech recognition task increases from 190.5 seconds to 191.6 seconds on Librispeech 100h dataset, as increases of only 1 second. Instead of computing higher-order polynomial terms, our GFSA approximates them, with only a small increase in runtime, which is not very significant.
@@
Table 26: Training time (seconds per epoch) on graph-level tasks
@@
Appendix S Inference Time Analysis
We report the inference time of various methods with GFSA in Tables 29 to 34.
@@
Table 32: Inference time on graph-level tasks
@@
Appendix T Results of the Strategy for Efficiency
In Tables 35 to 39, the results show that this strategy can reduce the increase in runtime and maintain performance compared to using GFSA for all layers.
@@
Appendix U GFSA in Linear Transformers
Table 40 shows the accuracy, runtime, and GPU usage results on Long Range Arena benchmark using ListOps and Image datasets.
Table 40: Comparison of accuracy (%), runtime ($s$ per 1,000 steps) and GPU usage (GB) on Long Range Arena benchmark
Method ListOps (2K) Image (4K)
Accuracy Runtime GPU usage Accuracy Runtime GPU usage
Transformer [77] 37.1 198.3 5.50 38.2 345.1 5.88
Transformer+GFSA 37.6 635.8 10.87 40.2 737.2 11.20
Linformer [81] 37.3 63.4 1.73 37.8 158.5 3.45
YOSO-E [97] 37.3 85.7 0.37 39.8 114.2 1.42
Efficient Attention [70] 36.9 49.2 0.57 40.2 121.1 1.14
Efficient Attention + GFSA 37.9 53.8 0.67 40.4 135.8 1.33